OpenStudy (dragon123mama):

Will award medal. i don't know how to solve this. What is the vertex form of the equation? y= -X^2 +12x -4

OpenStudy (dragon123mama):

@Mariamikayla

OpenStudy (mariamikayla):

the maximal point is obtain for a <0 and it's obtained for the equation---> -Delta/4a a= -1 <0 b=12 c= -4 Delta= b^2-4ac Delta= 12^2 -4 * (-1)* (-4)= 144 -16= 128 Vertex(f) = - delta/4a = -128/4*(-1) = -128/ -4= 32

OpenStudy (dragon123mama):

i still don't understand why you are using Delta. i never saw that used in the unit before...?

OpenStudy (mariamikayla):

because that's how i've learned :))

OpenStudy (dragon123mama):

f (x) = a(x - h)2 + k this is vertex form and your answer doesn't look like it is in that form .

OpenStudy (mariamikayla):

i'll try in that form, give me a moment :))

OpenStudy (dragon123mama):

kk

OpenStudy (mariamikayla):

f(x)=a(x-h)^2 + k h= -b/2a k= f(h) a= -1 b= 12 c= 4 h= -12/2*(-1) = -12/-2= 6 h=6 f(x)=ax^2+bx+c <=> f(h)=ah^2+bh+c k=f(h)=f(6)=a*6^2+b*6+c= 36a+6b+c k=36a+6b+c=36* (-1)+6*12+4= -36+72+4= -36+76= 40 k=40 f(x)= -a(x-h)^2+k f(x)= -(-1)(x-6)^2+40 f(x)= (x-6)^2 +40

OpenStudy (dragon123mama):

why are their 3 equations. which one is the answer.?

OpenStudy (dragon123mama):

wait nvm

OpenStudy (dumbcow):

@Mariamikayla , you have a typo c = -4 answer is: -(x-6)^2 +32

OpenStudy (mariamikayla):

ohhh....sorry...